OpenStudy (anonymous):

For y=x^2+2x-8, a)Find the the y-intercept and any x-intercepts b)Find the vertex c)Graph I'm having trouble with this problem because my vertex isn't the lowest point of the parabola and I don't see what I am doing wrong. My x-int.: (-4,0)(2,0) y-int.:(0,8) Vert: (-1, -5)

OpenStudy (helder_edwin):

if what you wrote if correct then \[ y=x^2+2x-8 \] \[ y+8+1=x^2+2x+1 \] \[ y+9=(x+1)^2 \] what are the coordinates for the vertx?